[Takafumi Ushiba, Yuta Michimura, Haoyu Wang]
After installing the whitening filter for POS PDs (IR s-pol and IR p-pol), see klog#21952, we did power measurements for the s-pol and p-pol with the current configurations for POP and POS. We didn't change any optics yet. The POP power is measured with CCD cameras following the same method in klog#21759. The power at POS is measured with the two DCPDs for s-pol (K1:LSC-POS_SPOL_DC_OUT16) and p-pol (K1:LSC-POS_PPOL_DC_OUT16). We found the POS s-pol and p-pol signals fluctuate quite a lot. We checked their spectrum and found this fluctuation comes from the peak at around 120Hz. So we take the average of the 10 second data for each signal for four times, and then do the average of the 4 data slices.
| POP SUM | POP EXP | POP SUM/EXP | POP_P SUM | POP_P EXP | POP_P SUM/EXP | POS s-pol | POS p-pol | |
| ITMX single | 4.9e6 | 750 | 6533.3 | 8.2e6 | 10000 | 820 | 1919±10 | 50.7±0.4 |
| Xarm locked | 4.8e6 | 750 | 6400 | 7.5e5 | 10000 | 75 | 1933±21 | 23.8±0.5 |
| ITMY single | 4.7e6 | 750 | 6266.7 | 4.6e6 | 10000 | 460 | 1785±5 | 342±1 |
| Yarm locked | 4.8e6 | 750 | 6400 | 6.5e5 | 10000 | 65 | 1830±10 | 139±5 |
After the measurements, we worried the gate valve in the tube may introduce clipping of the beam and we wanted to measure the mode matching ratio. We do this with the green laser with CARM and DARM locked. Then we add offset to the green CARM gradually and watch the TMS PDs for both arms. The mode mismatch is quite small (<1%). The misalignment (TEM01) is around 10% for both arms.